# Thermodynamics of Black Holes, far from Equilibrium

### v1

@@ Abstract
However, thanks to the BH uniqueness theorem [38, 39], one can recast the standard first law using quantities that are specified just at $H$. First note that the horizon areal radius $R$ (defined by $A=4\pi R^{2}$), and the horizon angular momentum $J_{H}=-\frac{1}{8\pi G}\oint\epsilon_{abcd}(\nabla^{a}\varphi^{b})\,{\rm d}S^{cd}$ (the Komar integral of $\mathring{\varphi}^{a}$ at $H$) are determined by fields at the horizon. Secondly, a Kerr black hole can be characterized by values of $(R,J_{H})$. Next, since the surface gravity $\mathring{\kappa}$ scales linearly with the horizon null normal, one can fix the rescaling freedom in $\mathring{k}^{a}$ by requiring that its surface gravity $\mathring{\kappa}$ –also determined by the horizon geometry– be given by $\mathring{\kappa}=\kappa_{{}_{\rm Kerr}}(R,J)$. As noted above, there is no rescaling freedom in $\mathring{\varphi}^{a}$. Therefore by setting $\mathring{\Omega}=\mathring{\Omega}_{{}_{\rm Kerr}}(R,J)$, one can recover the restriction of $\mathring{t}^{a}$ to $H$ via $\mathring{t}^{a}\,\hat{=}\,\mathring{k}^{a}-\mathring{\Omega}\mathring{\varphi}^{a}$. (Throughout $\hat{=}\,$ will stand for equality at the QLH under consideration.) Then, the three (Komar) charges evaluated at the horizon are given by $Q_{{}^{(\mathring{k})}}\,\hat{=}\,\frac{\mathring{\kappa}\,A}{4\pi G},\,Q_{(\mathring{\varphi})}\,\hat{=}\,-2J_{H}$  and  $Q_{(\mathring{t})}\,\hat{=}\,\frac{\mathring{\kappa}\,A}{4\pi G}+2\mathring{\Omega}J_{H}$. They satisfy the first law
@@
Recall first that the equilibrium state $\mathfrak{e}$ of a black hole is characterized by two numbers $(R,J_{H})$. Thus the space $\mathcal{E}$ of equilibrium states is 2-dimensional. By contrast, the space $\mathcal{N}$ of non-equilibrium states is infinite dimensional since a non-equilibrium state $\mathfrak{n}$ is characterized by the plethora of fields on MTSs $S$ of a DHS $\mathcal{H}$ (that evolve from one MTS to another). One of them, the 2-metric $\tilde{q}_{ab}$ on any $S$, determines its areal radius $R[S]$ of $S$. To obtain the analog of $J_{H}$ we will restrict ourselves to DHS $\mathcal{H}$ that (are either axisymmetric or) approach a Kerr IHS in the asymptotic future (or past). Then, each $S$ also carries an infinite set of multipoles that characterize its shape and spin structure [43, 21, 31]. Among them is the angular momentum dipole $\mathcal{J}_{\mathcal{H}}[S]\,\hat{=}\,-\frac{1}{8\pi G}\oint_{S}K_{ab}\varphi^{a}dS^{b}$ where $K_{ab}$ is the extrinsic curvature of $\mathcal{H}$, and $\varphi^{a}$ an invariantly constructed rotational vector field on $\mathcal{H}$, which is: (i) tangential to and divergence-free on every MTS $S$; and, (ii) coincides with the axial Killing field if the intrinsic metric $q_{ab}$ of $\mathcal{H}$ is axisymmetric. Thus, on the space $\mathcal{N}$ of non-equilibrium states we have again a pair of observables, $(R[S],\mathcal{J}_{\mathcal{H}}[S])$, that have the same physical interpretation as $(R,J_{H})$ have on $\mathcal{E}$. Interestingly, $\varphi^{a}$ can be naturally extended off $\mathcal{H}$ so that its Komar integral $Q_{(\varphi)}[S]$ equals $2\mathcal{J}_{\mathcal{H}}$, mirroring the relation $Q_{(\mathring{\varphi})}\,\hat{=}\,-2J_{H}$ on KHs.
@@
Therefore, there is a natural projection map $\Pi:\mathcal{N}\rightarrow\mathcal{E}$ such that $\Pi(\mathfrak{n})=\mathfrak{e}$ iff $(R,J_{H})=(R[S],\mathcal{J}_{\mathcal{H}}[S])$. Via pull-back, we can assign to any $\mathfrak{n}$, the intensive parameters $\mathring{\kappa},\,\mathring{\Omega}$ carried by $\mathfrak{e}$ (whose values are determined by $(R,J_{H})$). Thus, thanks to the BH uniqueness theorems, a key obstacle to extending thermodynamics to non-equilibrium situations is overcome: each non-equilibrium state $\mathfrak{n}$ has been assigned intensive parameters $\mathring{\kappa},\mathring{\Omega}$. Of course, $\mathring{\kappa},\,\mathring{\Omega}$ change in time on $\mathcal{H}$ since $(R[S],\mathcal{J}_{\mathcal{H}}[S])$ change from one MTS to another along $\mathcal{H}$, reflecting the fact that $\mathcal{H}$ represents a BH that is not in equilibrium.
@@
However, as remarked earlier, BHs are also more complicated! In thermodynamics, the notion of total energy of a system in a non-equilibrium state is well-defined. In general relativity, on the other hand, the notion of energy is tied to causal vector fields. On KHs, it was natural to use the horizon Killing field $\mathring{k}^{a}$ to define energy/charge $Q_{{}^{(\mathring{k})}}$ that appears on the left side of (1). What would be its analog on DHSs? Recall that DHSs carry a preferred null direction field $k^{a}$ (with $\theta_{(k)}\,\hat{=}\,0$) which, furthermore, aligns with $\mathring{k}^{a}$ as a DHS reaches equilibrium, becoming an IHS (as in Fig. 1). Hence a natural replacement of $\mathring{k}^{a}$ would be an appropriately scaled vector $\xi^{a}$ that is parallel to $k^{a}$. Now, it was shown in [16] that, thanks to the constraint equations on the DHS, there is an invariantly defined charge $Q_{(\xi)}[S]$ associated with any such $\xi^{a}$. Suppose the infalling flux ends at a MTS $\mathring{S}$ of the QLH, so that the DHS transitions to an IHS at $\mathring{S}$. Then, the DHS and IHS assignments of charges at $\mathring{S}$ would agree iff $\xi^{a}$ is chosen such that the DHS charge $Q_{(\xi)}[\mathring{S}]$ equals the IHS charge $Q_{{}^{(\mathring{k})}}\,\hat{=}\,\frac{\mathring{\kappa}\,A}{4\pi G}|_{\mathring{S}}$ for any MTS $\mathring{S}$. Surprisingly, this consistency condition selects a unique $\xi^{a}$ on the DHS!
@@
1st and 2nd laws on DHSs:   For definiteness, let us first focus on space-like DHSs. Then, $\hat{\tau}^{a}$ the unit normal to a DHS $\mathcal{H}$ is time-like, and $\hat{r}^{a}$ the unit normal within $\mathcal{H}$ to MTSs $S$ is space-like, and $k^{a}\,\hat{=}\,\frac{1}{\sqrt{2}}(\hat{\tau}^{a}+\hat{r}^{a})$ and $\underline{k}^{a}\,\hat{=}\,\frac{1}{\sqrt{2}}\,(\hat{\tau}^{a}-\hat{r}^{a})$ are the two null normals to $S$. As in the BCH analysis, we will use the 4 constraint equations, but now on the DHS $\mathcal{H}$ rather than on a partial Cauchy surface extending from the horizon to spatial infinity. Specifically, we will use combinations of the type $(G_{ab}-8\pi G\,T_{ab})\,\xi^{a}\hat{\tau}^{b}\,\hat{=}\,0$  of constraints, with $\xi^{a}\,\hat{=}\,\sqrt{2}\,|DR|f(R)\,k^{a}$ choosing functions $f(R)$ appropriately. (For motivation, see [16].)
On the DHS $\mathcal{H}$, $\mathcal{J}_{\mathcal{H}}[S]$ is a function of $R$, whence $\mathring{\kappa}$ and $\mathring{\Omega}$ are also functions of $R$, reflecting their dynamical nature. The unique choice of $\xi^{a}$ for which (the $\xi$-energy charge) $Q_{(\xi)}[S]\,\hat{=}\,Q_{(\mathring{k})}\equiv\frac{\mathring{\kappa}\,A}{4\pi G}$, as desired, is given by setting $f(R)\,\hat{=}\,\frac{{\rm d}}{{\rm d}R}\,(\frac{\mathring{\kappa}\,A}{2\pi G})$. Our constraint equation provides a balance law for this $Q_{(\xi)}$. Setting $\Delta\,[Q_{(\xi)}]\,\hat{=}\,Q_{(\xi)}[S_{1}]-Q_{(\xi)}[S_{2}]$, one has
@@
Here $\mathcal{H}_{1}^{2}$ is the portion of the DHS bounded by MTSs $S_{1}$ and $S_{2}$, and $\mathfrak{F}_{(\xi)}$ the flux of $\xi$-energy, with flux densities
$\mathfrak{f}_{(\xi)}^{\,\,\rm matt}\,\hat{=}\,T_{ab}\,\xi^{a}\hat{\tau}^{b}$   and,
@@
where $\sigma_{ab}$ is the shear of the null normal $k^{a}$, and $\zeta^{a}\,\hat{=}\,\,\tilde{q}^{ab}\,\hat{r}^{c}\nabla_{c}k_{b}$. The appearance of $|\zeta|^{2}$ seems surprising at first because it is absent in perturbative expression of the energy carried by gravitational waves across KHs. It arises because $\mathcal{H}$ is space-like rather than null: While there are only 2 true phase space degrees of freedom on null surfaces in general relativity, on a space-like surface there are 4, now captured in the fields $\sigma_{ab},\,\zeta^{a}$ tangential to MTSs, each with 2 free components [21]. They all contribute to the energy flux, as is usual in field theories.
@@
The second and third differences are intertwined. This becomes transparent in the infinitesimal form of the integral version, obtained by letting $S_{2}$ approach $S_{1}$ so that they are infinitesimally separated. Then our first law reduces to $\delta Q_{(t)}\,\,\hat{=}\,\,\delta[\frac{\mathring{\kappa}A}{4\pi G}]\,+\,2\delta[\mathring{\Omega}\mathcal{J}_{\mathcal{H}}]$, where $\delta$ denotes infinitesimal changes at $S_{1}$, along $\mathcal{H}$. Since every DHS defines a trajectory  $\mathfrak{e}(R)$  on the space $\mathcal{E}$ of equilibrium states, $S_{1}$ corresponds to a point $\mathfrak{e}_{1}$ in $\mathcal{E}$, and by construction the infinitesimal changes $\delta Q_{(t)},\,\delta[\frac{\mathring{\kappa}A}{4\pi G}]$ and $\delta[\mathring{\Omega}\mathcal{J}_{\mathcal{H}}]$ on $\mathcal{H}$ equal infinitesimal changes $\delta Q_{(\mathring{t})},\,\delta[\frac{\mathring{\kappa}A}{4\pi G}]$ and $\delta[\mathring{\Omega}J_{H}]$ at $\mathfrak{e}_{1}$, along $\mathfrak{e}(R)$. Now, thanks to the Smarr relations on the KH, we have an identity on $\mathcal{E}$: $\frac{A\,\delta\mathring{\kappa}}{4\pi G}+2J_{H}\delta\mathring{\Omega}\,\hat{=}\,-\big(\frac{\mathring{\kappa}\,\delta A}{8\pi G}+\mathring{\Omega}\delta J\big)$. By pulling it back to $\mathcal{H}$ via $\Pi$, the infinitesimal version of the DHS first law can be rewritten as
@@
which has exactly the same form as the first law (1) on KHs. But on the DHS the infinitesimal changes are caused by physical fluxes into $\mathcal{H}$ (near $S_{1}$); it is not a passive change from one equilibrium state to a nearby one. The rewriting procedure cannot be repeated to the finite version (4) of the first law because $\mathring{\kappa},\mathring{\Omega}$ are dynamical on $\mathcal{H}$ and therefore vary along the trajectory  $\mathfrak{e}(R)$  in $\mathcal{E}$.
This concludes our discussion of the generalization of the first law to dynamical BHs in general relativity. The second law for DHSs results from the combination $(G_{ab}-8\pi G\,T_{ab}){\underline{\xi}}^{a}\hat{\tau}^{b}\,\hat{=}\,0$, where ${\underline{\xi}}^{a}:=\sqrt{2}|DR|(4\pi R)\,k^{a}$. This constraint implies [16]:
@@
Discussion:   Space-times representing dynamical BHs do not admit KHs. They do admit EHs, but the growth in the area of EHs is unrelated to local physical processes. Therefore, the EH area cannot be a viable measure of the physical entropy in non-equilibrium situations. DHSs share some of the attractive properties with EHs –e.g., their area grows in classical GR, and they can merge but cannot bifurcate [22]. But DHSs are defined quasi-locally and are free from teleology. Therefore, they have been widely used to represent black hole far from equilibrium both und classical and quantum gravity. In this Letter we showed that they are also well-suited for extending BH thermodynamics to non-equilibrium situations. In particular, the expressions of our first and second law, and the fact that the time dependence of the area of MTSs of a DHS is governed by local physical processes, naturally lead one to identify the area of MTSs of a DHS with BH entropy in non-equilibrium situations. Finally, note that our analysis used structures that are directly available in the physical space-time. For example, one does not need bifurcate horizons that are often used, but do not exist in physical space-times of dynamical BHs beyond perturbation theory.
@@
2. In our discussion of the first and second law, black holes that are not in equilibrium were represented by DHSs. However, generically, in each of its dynamical phases, the BH can admit a number of ‘interweaving’ DHSs [22]. A key point is that each DHS provides us with a detailed, self-consistent description of the evolution of the BH and all our thermodynamical considerations hold for each of them. When the BH reaches equilibrium, represented by an IH, these DHSs asymptote to that IH and their thermodynamic parameters agree with those of the IH. Thus, there is consistency between the dynamical and equilibrium phases.
3. Recently, phase space methods were used to show that, already when first order dynamical perturbations are included, the entropy has to be associated not with the area of the EH, but with areas of a family of MTSs that lie inside the EH [44, 45]. The world-tube of these MTSs can be thought of as a perturbative DHS and qualitative features of the perturbative analysis can be read off from our non-perturbative expressions. However, conceptually the two frameworks are rather different. Similarities and differences will be discussed in [46]. We will also discuss charges associated with gauge fields and time-like DHSs there. However, so far our analysis has been focused on QLHs in general relativity. QLHs do exist in other gravitational theories and some of their key properties follow simply from the geometric Gauss-Codazzi equations [21]. It would be interesting to investigate whether ideas underlying our analysis can be extended to other metric theories of gravity.

### v2

@@ Abstract
Fortunately, thanks to the BH uniqueness theorem [35, 26], one can recast the standard first law using quantities that are specified just at $H$ and then extend it to dynamical situations. First note that the horizon areal radius $R$ (defined by $A=4\pi R^{2}$), and the horizon angular momentum $J_{H}=\frac{1}{16\pi G}\oint\epsilon_{abcd}(\nabla^{a}\varphi^{b})\,{\rm d}S^{cd}$ ($\frac{1}{2}$ the Komar integral of $\mathring{\varphi}^{a}$ at $H$) are determined by fields at the horizon. Second, a Kerr black hole is fully characterized by values of $(R,J_{H})$. Third, since the surface gravity $\mathring{\kappa}$ scales linearly with the horizon null normal, one can fix the rescaling freedom in $\mathring{k}^{a}$ by requiring that its surface gravity $\mathring{\kappa}$ –also determined by the horizon geometry– be given by $\mathring{\kappa}=\kappa_{{}_{\rm Kerr}}(R,J)$. As noted above, there is no rescaling freedom in $\mathring{\varphi}^{a}$. Therefore by setting $\mathring{\Omega}=\mathring{\Omega}_{{}_{\rm Kerr}}(R,J)$, one can recover the restriction of $\mathring{t}^{a}$ to $H$ via $\mathring{t}^{a}\,\hat{=}\,\mathring{k}^{a}-\mathring{\Omega}\mathring{\varphi}^{a}$. (Throughout $\hat{=}\,$ will stand for equality at the QLH under consideration.) Then, the three (Komar) charges evaluated at the horizon are given by $Q_{{}^{(\mathring{k})}}\,\hat{=}\,\frac{\mathring{\kappa}\,A}{4\pi G},\,Q_{(\mathring{\varphi})}\,\hat{=}\,-2J_{H}$  and  $Q_{(\mathring{t})}\,\hat{=}\,\frac{\mathring{\kappa}\,A}{4\pi G}+2\mathring{\Omega}J_{H}$. They satisfy the first law
@@
Recall first that the equilibrium state $\mathfrak{e}$ of a black hole is characterized by two numbers $(R,J_{H})$. Thus the space $\mathcal{E}$ of equilibrium states is 2-dimensional. By contrast, the space $\mathcal{N}$ of non-equilibrium states $\mathfrak{n}$ is infinite dimensional since each $\mathfrak{n}$ is characterized by the plethora of fields on MTSs $S$ of a DHS $\mathcal{H}$ (that evolve from one MTS to another). One of them, the 2-metric $\tilde{q}_{ab}$ on any $S$, determines its areal radius $R[S]$ of $S$. To obtain the analog of $J_{H}$ we will restrict ourselves to DHS $\mathcal{H}$ that (are either axisymmetric or) approach a Kerr IHS in the asymptotic future (or past). Then, each $S$ also carries an infinite set of multipoles that characterize its shape and spin structure [7, 13, 25]. Among them is the angular momentum dipole $\mathcal{J}_{\mathcal{H}}[S]\,\hat{=}\,-\frac{1}{8\pi G}\oint_{S}K_{ab}\varphi^{a}dS^{b}$ where $K_{ab}$ is the extrinsic curvature of $\mathcal{H}$, and $\varphi^{a}$ an invariantly constructed rotational vector field on $\mathcal{H}$, which is: (i) tangential to and divergence-free on every MTS $S$; and, (ii) coincides with the axial Killing field if the intrinsic metric $q_{ab}$ of $\mathcal{H}$ is axisymmetric. Thus, on the space $\mathcal{N}$ of non-equilibrium states we have again a pair of observables, $(R[S],\mathcal{J}_{\mathcal{H}}[S])$, that have the same physical interpretation as $(R,J_{H})$ have on $\mathcal{E}$. (Interestingly, $\varphi^{a}$ can be naturally extended off $\mathcal{H}$ so that its Komar integral $Q_{(\varphi)}[S]$ equals $2\mathcal{J}_{\mathcal{H}}$, mirroring the relation $Q_{(\mathring{\varphi})}\,\hat{=}\,-2J_{H}$ on KHs.)
@@
Therefore, there is a natural projection map $\Pi:\mathcal{N}\rightarrow\mathcal{E}$ such that $\Pi(\mathfrak{n})=\mathfrak{e}$ iff $(R,J_{H})=(R[S],\mathcal{J}_{\mathcal{H}}[S])$. Via pull-back, we can assign to any $\mathfrak{n}$, the intensive parameters $(\mathring{\kappa},\,\mathring{\Omega})$ carried by $\mathfrak{e}$. Thus, thanks to the BH uniqueness theorems, a key obstacle to extending thermodynamics to non-equilibrium situations is overcome: each non-equilibrium state $\mathfrak{n}$ has been assigned intensive parameters. Of course, on $\mathcal{H}$  $(\mathring{\kappa},\,\mathring{\Omega})$ change in time since $(R[S],\mathcal{J}_{\mathcal{H}}[S])$ change from one MTS to another, reflecting the fact that $\mathcal{H}$ represents a dynamical BH.
@@
However, as remarked earlier, BHs are also more complicated! In thermodynamics, the notion of total energy of a system in a non-equilibrium state is well-defined. In general relativity, on the other hand, the notion of energy is tied to causal vector fields. On KHs, it was natural to use the horizon Killing field $\mathring{k}^{a}$ to define energy/charge $Q_{{}^{(\mathring{k})}}$ that appears on the right side of (1). What would be its analog on DHSs? Recall that DHSs carry a preferred null direction field $k^{a}$ (with $\theta_{(k)}\,\hat{=}\,0$) which, furthermore, aligns with $\mathring{k}^{a}$ as a DHS reaches equilibrium, becoming an IHS (as in Fig. 1). Hence a natural replacement of $\mathring{k}^{a}$ would be an appropriately scaled vector $\xi^{a}$ that is parallel to $k^{a}$. Now, it was shown in [11] that, thanks to the constraint equations on the DHS, there is an invariantly defined charge $Q_{(\xi)}[S]$ associated with any such $\xi^{a}$. Suppose the infalling flux ends at a MTS $\mathring{S}$ of the QLH, so that the DHS transitions to an IHS at $\mathring{S}$. Then, the DHS and IHS assignments of charges at $\mathring{S}$ would agree if $\xi^{a}$ is chosen such that the DHS charge $Q_{(\xi)}[\mathring{S}]$ equals the IHS charge $Q_{{}^{(\mathring{k})}}\,\hat{=}\,\frac{\mathring{\kappa}\,A}{4\pi G}|_{\mathring{S}}$ for any MTS $\mathring{S}$. Surprisingly, this consistency condition selects a unique $\xi^{a}$ on the DHS!
@@
1st and 2nd laws on DHSs:   For definiteness, let us first focus on space-like DHSs. Then, $\hat{\tau}^{a}$ the unit normal to a DHS $\mathcal{H}$ is time-like; $\hat{r}^{a}$ the unit normal within $\mathcal{H}$ to MTSs $S$ is space-like; and $k^{a}\,\hat{=}\,\frac{1}{\sqrt{2}}(\hat{\tau}^{a}+\hat{r}^{a})$ and $\underline{k}^{a}\,\hat{=}\,\frac{1}{\sqrt{2}}\,(\hat{\tau}^{a}-\hat{r}^{a})$ are the two null normals to $S$. As in the BCH analysis, we will use the 4 constraint equations, but now on the DHS $\mathcal{H}$ rather than on a partial Cauchy surface extending from the horizon to spatial infinity. Specifically, we will use combinations of the type $(G_{ab}-8\pi G\,T_{ab})\,\xi^{a}\hat{\tau}^{b}\,\hat{=}\,0$  of constraints, with $\xi^{a}\,\hat{=}\,\sqrt{2}\,|DR|f(R)\,k^{a}$ choosing functions $f(R)$ appropriately. (For motivation, see [11].)
On the DHS $\mathcal{H}$, $\mathcal{J}_{\mathcal{H}}[S]$ is a function of $R$, whence $\mathring{\kappa}$ and $\mathring{\Omega}$ are also functions of $R$, reflecting their dynamical nature. The unique choice of $\xi^{a}$ for which (the $\xi$-energy charge) $Q_{(\xi)}[S]\,\,\hat{=}\,Q_{(\mathring{k})}\equiv\frac{\mathring{\kappa}\,A}{4\pi G}$, is given by setting $f(R)\,\hat{=}\,\frac{{\rm d}}{{\rm d}R}\,(\frac{\mathring{\kappa}\,A}{2\pi G})$. Our constraint equation provides a balance law for this $Q_{(\xi)}$. Setting $\Delta\,[Q_{(\xi)}]\,\,\hat{=}\,Q_{(\xi)}[S_{2}]-Q_{(\xi)}[S_{1}]$, one has
@@
Here $\mathcal{H}_{1}^{2}$ is the portion of the DHS bounded by MTSs $S_{1}$ and $S_{2}$, and $\mathfrak{F}_{(\xi)}$ the flux of $\xi$-energy, with flux densities for matter fields and gravitational waves
$\mathfrak{f}_{(\xi)}^{\,\,\rm matt}\,\hat{=}\,T_{ab}\,\xi^{a}\hat{\tau}^{b}$   and
@@
where $\sigma_{ab}$ is the shear of the null normal $k^{a}$, and $\zeta^{a}\,\hat{=}\,\,\tilde{q}^{ab}\,\hat{r}^{c}\nabla_{c}k_{b}$. The appearance of $|\zeta|^{2}$ seems surprising at first because it is absent in perturbative expressions of the energy carried by gravitational waves across KHs. It arises because $\mathcal{H}$ is space-like rather than null: While there are only 2 true phase space degrees of freedom on null surfaces in general relativity, on a space-like surface there are 4, now captured in the fields $(\sigma_{ab},\,\zeta^{a})$ tangential to MTSs, each with 2 free components [13]. They all contribute to the energy flux, as is usual in field theories.
@@
The second and third differences are intertwined. This becomes transparent in the infinitesimal form of the integral version, obtained by letting $S_{2}$ approach $S_{1}$ so that they are infinitesimally separated. Then our first law reduces to $\delta Q_{(t)}\,\,\hat{=}\,\,\delta[\frac{\mathring{\kappa}A}{4\pi G}]\,+\,2\delta[\mathring{\Omega}\mathcal{J}_{\mathcal{H}}]$, where $\delta$ denotes infinitesimal changes at $S_{1}$, along $\mathcal{H}$. Since every DHS defines a trajectory  $\mathfrak{e}(R)$  on the space $\mathcal{E}$ of equilibrium states, $S_{1}$ corresponds to a point $\mathfrak{e}_{1}$ in $\mathcal{E}$. Thanks to the properties of the projection $\Pi$, the infinitesimal changes $\delta Q_{(t)},\,\delta[\frac{\mathring{\kappa}A}{4\pi G}]$ and $\delta[\mathring{\Omega}\mathcal{J}_{\mathcal{H}}]$ on $\mathcal{H}$ equal infinitesimal changes $\delta Q_{(\mathring{t})},\,\delta[\frac{\mathring{\kappa}A}{4\pi G}]$ and $\delta[\mathring{\Omega}J_{H}]$ at $\mathfrak{e}_{1}$, along $\mathfrak{e}(R)$. Now, the Smarr relations on the KH imply an identity on $\mathcal{E}$: $\frac{A\,\delta\mathring{\kappa}}{4\pi G}+2J_{H}\delta\mathring{\Omega}\,\hat{=}\,-\big(\frac{\mathring{\kappa}\,\delta A}{8\pi G}+\mathring{\Omega}\delta J\big)$. By pulling it back to $\mathcal{H}$ via $\Pi$, the infinitesimal version of the DHS first law can be rewritten as
@@
which has exactly the same form as the first law (1) on KHs. (But on the DHS the infinitesimal changes are caused by physical fluxes into $\mathcal{H}$ (near $S_{1}$); it is not a passive change from one equilibrium state to a nearby one.) The finite version (4) of the first law cannot be rewritten in this way: since $(\mathring{\kappa},\mathring{\Omega})$ are dynamical on $\mathcal{H}$, they vary along the trajectory  $\mathfrak{e}(R)$  in $\mathcal{E}$, whence the identity we used does not hold for finite variations.
This concludes our discussion of the generalization of the first law to dynamical BHs in general relativity. The second law for DHSs results from the combination $(G_{ab}-8\pi G\,T_{ab}){\underline{\xi}}^{a}\hat{\tau}^{b}\,\hat{=}\,0$, where ${\underline{\xi}}^{a}:=\sqrt{2}|DR|(4\pi R)\,k^{a}$. This constraint implies:
@@
Discussion:  Space-times representing dynamical BHs do not admit KHs. They do admit EHs, but the growth in the area of EHs is unrelated to local physical processes. Therefore, the EH area cannot be a viable measure of the physical entropy in non-equilibrium situations. DHSs share some attractive properties with EHs –e.g., their area grows in classical GR, and they can merge but cannot bifurcate [9]. Furthermore, DHSs are defined quasi-locally and are free from teleology. Therefore, they have been widely used to represent black holes far from equilibrium both in classical and quantum gravity. In this Letter we showed that they are also well-suited for extending BH thermodynamics to non-equilibrium situations. In particular, the expressions of our first and second law, and the fact that the time dependence of the area of MTSs of a DHS is governed by local physical processes, naturally lead one to identify the area of MTSs of a DHS with BH entropy in non-equilibrium situations. Finally, note that our analysis used structures that are directly available in the physical space-time. For example, one does not need bifurcate horizons that are often used, but do not exist in physical space-times of dynamical BHs beyond perturbation theory.
@@
2. We used DHSs to represent BHs that are not in equilibrium. Generically, in each of its dynamical phases, the BH can admit a number of ‘interweaving’ DHSs $\mathcal{H}$ [9]. A key point is that each $\mathcal{H}$ provides us with a detailed, self-consistent description of the evolution of the BH and all our thermodynamical considerations also hold for each $\mathcal{H}$. When the BH reaches equilibrium, represented by an IHS, all these DHSs asymptote to that IHS and their thermodynamic parameters agree with those of the IHS [7]. Thus, there is consistency between the dynamical and equilibrium phases.
3. Recently it was shown that, already when first order, dynamical perturbations are included around stationary BHs, entropy is given, not with the area of the EH, but with areas of a family of MTSs that lie inside the EH [36, 46]. The world-tube of these MTSs can be thought of as a perturbative DHS and qualitative features of the perturbative analysis can be read off from our non-perturbative expressions. However, conceptually the two frameworks are rather different. Similarities and differences will be discussed in [16]. We will also discuss charges associated with gauge fields and time-like DHSs in quantum gravity there. However, so far our analysis has been focused on QLHs in general relativity. QLHs do exist in other gravitational theories and some of their key properties follow simply from the geometric Gauss-Codazzi equations [13]. It would be interesting to investigate whether ideas underlying our analysis can be extended to other metric theories of gravity.